Thanks in advance!

Part A Find the place(s) where the electrc field is zero Express your answer using two decimal places. f there is more than one answer, separate them by commas 255 8 Submt Provide Feedback
Answer #1

Let's assume a point O, at x distance from q1 where electric field is 0 (point x lie on -x axis).

kq1 / x^2 - kq2 / x2^2 = 0

where, x2 = distance of point O from q2

x2 = x + (2.55 - 1) = x + 1.55

kq1 / x^2 - kq2 / (x + 1.55)^2 = 0

1.65*10^(-6) / x^2 = 2.20*10^(-6) / (x + 1.55)^2

take root both the sides,

1.284 / x = 1.483 / (x + 1.55)

1.284x + 1.991 = 1.483x

0.198x = 1.991

x = 10.01 m

distance of point O from origin,

x = 10.018 - 1

x = 9.018 m (at -x axis)